### v1

# RedDebate: Safer Responses through Multi-Agent Red Teaming Debates

Abstract
We propose RedDebate, a novel multi-agent debate framework that leverages adversarial argumentation among Large Language Models (LLMs) to proactively identify and mitigate their own unsafe behaviours. Existing AI safety methods often depend heavily on costly human evaluations or isolated single-model assessment, both subject to scalability constraints and oversight risks. RedDebate instead embraces collaborative disagreement, enabling multiple LLMs to critically examine one another’s reasoning, and systematically uncovering unsafe blind spots through automated red-teaming, and iteratively improve their responses. We further integrate distinct types of long-term memory that retain learned safety insights from debate interactions. Evaluating on established safety benchmarks such as HarmBench, we demonstrate the proposed method’s effectiveness. Debate alone can reduce unsafe behaviours by 17.7%, and when combined with long-term memory modules, achieves reductions exceeding 23.5%. To our knowledge, RedDebate constitutes the first fully automated framework that combines multi-agent debates with red-teaming to progressively enhance AI safety without direct human intervention. [^1] https://github.com/aliasad059/RedDebate
RedDebate: Safer Responses through Multi-Agent Red Teaming Debates
Ali Asad, Stephen Obadinma, Radin Shayanfar, Xiaodan Zhu Department of Electrical and Computer Engineering & Ingenuity Labs Research Institute Queen’s University, Kingston, Canada {ali.asad, 16sco, radin.shayanfar, xiaodan.zhu}@queensu.ca
1 Introduction
“It is better to change an opinion than to persist in a wrong one.” —Socrates
Humans have long employed debate and dialogue as powerful methods to refine their understanding and correct misconceptions, a tradition deeply rooted in the Socratic method, or Elenchus Seeskin (1987); Paul and Elder (2008). Recently, advances in Artificial Intelligence (AI), specifically in Large Language Models (LLMs) have equipped machines with unprecedented capabilities for natural language communication. These advances include enhanced understanding, reasoning, and generating of language grounded in extensive accumulated knowledge (Minaee et al., 2024; OpenAI, 2024). These foundational skills provide an opportunity to explore debate as a means to leverage the inherent strengths of these models (Bench-Capon and Dunne, 2007). Specifically, debate can facilitate examining complex ideas, evaluating evidence, and critically analyzing competing viewpoints. As a result, there is growing interest in using debate to enhance and harness the strengths of LLMs (Irving et al., 2018; Parrish et al., 2022; Du et al., 2024).
LLM-based debate explicitly leverages the diversity of perspectives present within different LLMs, allowing these highly capable language models to contribute unique insights within structured dialogues (Liang et al., 2024). Therefore, this richer collaborative process has the potential to enhance the depth, factuality, and robustness of model-generated reasoning (Du et al., 2024; Khan et al., 2024). Indeed, debate has been shown to effectively surface flaws in reasoning, proving particularly valuable for enabling human judges or comparatively less capable AI systems to detect incorrect or potentially unsafe reasoning from superhuman-level AI (Irving et al., 2018; Khan et al., 2024). Thus, debating represents a promising approach to address key safety concerns within AI systems by effectively tapping into the advanced capabilities of modern LLMs and fostering collective reasoning.
Figure 1: RedDebate framework overview. Agents debate a red-teaming prompt over multiple rounds, refining responses through peer interaction. An evaluator performs a further safety check, and key safety insights are then generated and stored in memory for future use, enabling continuous safety improvement.
One prominent challenge in the AI safety domains is the detection and mitigation of unsafe AI responses—commonly addressed through red-teaming methods. These unsafe responses can be characterized as arising from the AI system’s misconceptions regarding alignment with human values. Traditional approaches to red-teaming and alignment heavily depend on human annotations and evaluations, facing challenges in scalability, due to critical oversights, and incur significant costs, as LLMs continue to grow in complexity and size (Bowman et al., 2022; Perez et al., 2022; Bai et al., 2022c). AI-driven supervision methods present an alternative by automating these processes through the enhanced reasoning capabilities of LLMs. Most of these approaches rely on single-agent supervision, such as constitutional AI using self-criticism under predefined constitutions (Bai et al., 2022c) and self-reflection (Shinn et al., 2023). The key issue with these methods is that an agent might not realize its own mistakes or knowledge gaps, making it difficult to reflect and correct itself effectively (Liang et al., 2024).
To address these limitations, we propose RedDebate, a novel multi-agent debate framework designed to enhance AI safety by enabling collaborative, adversarial exploration of unsafe reasoning, fostering deeper insights unattainable through individual self-assessment alone. Our framework combines the collective reasoning strengths inherent in debate, including diversity of perspectives and mutual critical assessment, along with systematic vulnerability detection through automated red-teaming. With this objective in mind, we explore the central research question: Can a set of LLMs collaboratively identify and refine their unsafe behaviours through debate?
As illustrated in Figure 1, RedDebate facilitates mutual reflection among agents on the safety of their own and each other’s responses. An evaluator further assesses response safety, providing feedback to guide further improvement, which agents retain as learned insights within their memory.
Our work makes three key contributions:
  • To our knowledge, we present the first fully automated learning framework that combines debate and red-teaming, enabling agents to collaboratively identify and refine unsafe responses without human intervention.
  • We introduce variant types of long-term memory tested across diverse debate scenarios, showing that memory significantly enhances agent performance via dynamic feedback retention and retrieval, while debates expose distinct vulnerabilities.
  • We demonstrate that LLMs effectively learn from debate and memory to reduce unsafe responses, achieving up to 17.7% reduction with debate alone and over 23.5% when combined with long-term memory on HarmBench.
2 Related Work
Multi-Agent Debate (MAD)
improves reasoning and factual accuracy by having multiple LLMs interact, each offering different perspectives. Prior work shows that debate outperforms single-agent methods like self-reflection or ensembling (Smit et al., 2024; Chan et al., 2024; Du et al., 2024; Liang et al., 2024). In early foundational work, Irving et al. (2018) introduced debate for superhuman AI alignment through self-play. Then, Khan et al. (2024) showed debate helps weaker judges evaluate stronger models. However, none of these works employ MAD as a red-teaming strategy combined with learning for safety refinement as we do.
Red Teaming
exposes unsafe model behaviour via adversarial prompts. While early work relied on human-written tests, recent methods automate attack generation and evaluation using LLMs (Perez et al., 2022), with extensions for broader coverage (Hong et al., 2024; Casper et al., 2023). Our work advances this by creating a fully automated red-teaming pipeline where LLM agents generate, critique, and evaluate responses.
Learning From Feedback
Feedback, either human or model-generated, helps steer LLM behaviour. RLHF finetunes models based on human preferences (OpenAI et al., 2024; Bai et al., 2022a), while Bai et al. (2022b) and Shinn et al. (2023) use self-generated feedback for response improvement via predefined rules or textual feedback derived in subsequent trials. We extend this with peer critique in debate to produce richer feedback, combined with long-term memory for enhanced retention.
Agent Memory
To overcome context limitations, LLM agents benefit from means of storing and accessing information they have previously learned across interactions (Zhang et al., 2024). Textual memory stores knowledge in natural language and retrieves it via different methods, among them vector search (Hu et al., 2023; Zhong et al., 2024), while parametric memory updates model weights through fine-tuning (Xiong et al., 2023). We are the first to use and integrate both memory types in a debate-based safety framework.
Guardrailing
allows determining if and how some actions could be enforced to increase safety in a system (Dong et al., 2024b; Rebedea et al., 2023; Guardrails AI, ). NeMo Guardrails (Rebedea et al., 2023) uses Colang [^2] https://docs.nvidia.com/nemo/guardrails/colang_2/overview.html to define safe and highly flexible conversational flows. Recently, the canonical form extraction of automatic guardrails (Sreedhar et al., 2024) has garnered interest and shown promising results in task-oriented dialogue systems. We build on this by introducing guardrails as a form of long-term memory for safety learning. Please refer to Appendix A for a more detailed overview of related works for each area.
3 Methodology
We introduce RedDebate, a multi-agent framework where agents collaboratively tackle red-teaming prompts and learn from failures.
Figure 2: RedDebate Framework
3.1 RedDebate Framework
As illustrated in Figure 2, the process begins by selecting a red-teaming prompt $\mathcal{X}\in\mathcal{P}$ from a set of adversarial prompts $\mathcal{P}$, which serves as the central topic for the debate. A set of $N$ debater agents $\mathcal{D}=\{\mathrm{D_{1}},\mathrm{D_{2}},\ldots,\mathrm{D_{N}}\}$ simultaneously generate responses at each round $t$, as defined in Equation 1,
$r_{n}^{(t)}=\mathrm{D_{n}}(\mathcal{X},\mathcal{M}_{\text{S}},\mathcal{M}_{\text{L}})$ (1)
where $\mathcal{M}_{\text{S}}$ and $\mathcal{M}_{\text{L}}$ denote the shared short-term and long-term memory, respectively. At each round $t$, the set of agents responses $\mathcal{R}^{(t)}=\{r_{n}^{(t)}\}_{n=1}^{N}$ will be stored in $\mathcal{M}_{\text{S}}$. This enables each agent to access not only its own previous answers but also those of other agents. In subsequent rounds, agents use this memory to critique others’ responses, refine their previous statements, or offer feedback to peers. After a fixed number of rounds of debate $T$, a separate evaluator agent $\mathrm{E}$ assesses the safety of all responses =$\mathcal{R}=\{r_{n}^{(t)}\}_{n=1,t=1}^{N,T}$ generated throughout the debate and produces binary labels $\mathcal{Y}=\{y_{n}^{(t)}\}_{n=1,t=1}^{N,T}$, as described by Equation 2, where each label $y_{n}^{(t)}\in\{0,1\}$ indicates whether the response is safe.
$\mathcal{Y}=\mathrm{E}(\mathcal{R})$ (2)
If any $y_{n}^{(t)}=0$, indicating that at least one unsafe response was produced during the debate, a feedback generator $\mathrm{F}$ receives the full debate history $\mathcal{R}$ and corresponding safety labels $\mathcal{Y}$, and produces a textual explanation $\phi$ highlighting the identified flaws—similar in spirit to the reflection mechanism in Shinn et al. (2023), though extended here to a multi-agent safety setting (Equation 3).
$\phi=\mathrm{F}(\mathcal{R},\mathcal{Y})$ (3)
The resulting feedback $\phi$, which represents a distilled safety insight, is stored in the shared long-term memory $\mathcal{M}_{\text{L}}$ accessible to all agents in future debates. This memory acts as a repository of lessons learned from previous unsafe behaviours, enabling agents to improve over time, independent of human supervision.
Importantly, the evaluator $\mathrm{E}$ is kept separate from the debaters $\mathcal{D}$ to provide an independent assessment. This is crucial because when all agents share a flawed belief and respond unsafely, they may be unable to correct each other. The evaluator acts as an external signal, flagging such coordinated failure and triggering feedback generation. Additionally, even when all agents respond safely, there remains a concern that the evaluator may have overlooked subtle unsafe content. However, since multiple agents produce diverse responses, there is a greater chance that at least one will flag or contradict a potentially unsafe answer, prompting further scrutiny. In this way, the debate mechanism enhances the robustness of safety evaluation by surfacing disagreements and divergent reasoning paths Chan et al. (2024); Chern et al. (2024). Algorithm 1 summarizes the RedDebate process.
1
Input: Prompt $\mathcal{X}$, debaters $\mathcal{D}=\{\mathrm{D_{1}},\dots,\mathrm{D_{N}}\}$, evaluator $\mathrm{E}$, feedback generator $\mathrm{F}$, short-term memory $\mathcal{M}_{\text{S}}$, long-term memory $\mathcal{M}_{\text{L}}$, rounds $T$
Output: Debate history $\mathcal{R}$
2
$\mathcal{M}_{\text{S}}\leftarrow\emptyset$
  // Initialize STM
$\mathcal{R}\leftarrow[]$
  // Initialize Debate History
3
4for $t\leftarrow 1$ to $T$ do
5       $\mathcal{R}^{(t)}\leftarrow\{\}$
6
7      for $n\leftarrow 1$ to $N$ do
             $r_{n}^{(t)}\leftarrow\mathrm{D_{n}}(\mathcal{X},\mathcal{M}_{\text{S}},\mathcal{M}_{\text{L}})$
              // Debate
8             $\mathcal{R}^{(t)}\leftarrow\mathcal{R}^{(t)}\cup\{r_{n}^{(t)}\}$
9
10       end for
11
      $\mathcal{M}_{\text{S}}\leftarrow\mathcal{M}_{\text{S}}\cup\mathcal{R}^{(t)}$
        // Update STM
       $\mathcal{R}\leftarrow\mathcal{R}\cup\{\mathcal{R}^{(t)}\}$
        // Append to History
12
13 end for
14
$\mathcal{Y}\leftarrow\mathrm{E}(\mathcal{R})$
  // Evaluate History
15
16if $\exists\,y_{n}^{(t)}=0$ in $\mathcal{Y}$ then
       $\phi\leftarrow\mathrm{F}(\mathcal{R},\mathcal{Y})$
        // Generate Feedback
       $\mathcal{M}_{\text{L}}\leftarrow\mathcal{M}_{\text{L}}\cup\{\phi\}$
        // Update LTM
17
18 end if
19
return $\mathcal{R}$
  // Return Debate History
Algorithm 1 Multi-Agent Debate with Memory
3.2 Debate Strategies
Exploring diverse debate strategies is essential (Chan et al., 2024; Smit et al., 2024), as different agent roles and communication styles can elicit varying perspectives and reasoning paths. We explore the following scenarios, each specifically designed for safety.
Peer Refinement Debate (PReD)
We initially introduce PReD, a simple debate strategy in which multiple peer agents with identical roles respond in parallel to a red-teaming prompt. This approach enables agents to collaboratively refine their potentially unsafe behaviours, following the same procedure as outlined in Algorithm 1.
Devil–Angel Refinement Debate (DAReD)
In PReD, often responses generated by debating agents can overlap or represent similar reasoning, reducing the diversity and critical scope of the discussion. To address this limitation, we explore the introduction of auxiliary agents that intentionally diversify the debate landscape. These agents enhance robustness by exposing the debating agent to explicitly contrasting perspectives, prompting a more critical reassessment of the original response.
To achieve our goal—and inspired by Liang et al. (2024) who promote perspective diversification, though in a different context—we integrate two auxiliary agents in the safety setting: the Devil agent ($\mathrm{D}^{\ominus}$) and the Angel agent ($\mathrm{D}^{\oplus}$). When a debating agent produces a round-$t$ response $r_{n}^{(t)}$, the Devil agent generates a rejection $\delta_{n}^{(t)}=\mathrm{D}^{\ominus}(r_{n}^{(t)})$, which critically challenges the response, acting as a skeptical adversary independent of the actual safety status. Conversely, the Angel agent provides supportive reinforcement $\alpha_{n}^{(t)}=\mathrm{D}^{\oplus}(r_{n}^{(t)})$, explicitly affirming and encouraging the original response. With these explicitly contrasting viewpoints—one inherently skeptical (possibly unsafe) and the other explicitly supportive (likely perceived as safe)—the agent is compelled to critically reassess its reasoning. This opposing feedback is anticipated to facilitate self-refinement and correction of potential unsafe responses. If the agent does not adequately reflect or becomes misled by persuasive auxiliary input, such inadequate responses are more reliably identified and flagged by the evaluator $\mathrm{E}$ and feedback generator $\mathrm{F}$ for further corrective intervention.
Socratic Refinement Debate (SReD)
We observe that agents often do not proactively provide feedback or refine their responses based on peer contributions unless explicitly triggered or challenged by others. Additionally, assessing the depth and stability of an agent’s conviction in its responses—whether these are firmly held or merely superficial positions subject to change—proved crucial for effective deliberation. To address these concerns, we introduce a novel scenario incorporating an auxiliary Socratic agent $\mathrm{D^{S}}$ alongside the basic debaters $\mathcal{D}$. Within this setting, the Socratic agent assumes a distinct questioning role, critically examining the responses $\mathcal{R}^{(t)}$ provided at debate round $t$. Its primary objective is to uncover implicit assumptions, highlight reasoning gaps, and promote deeper reflection by actively requesting evidence, illustrative examples, or clarification as needed. Inspired by traditional Socratic dialogues Paul and Elder (2008); Chang (2023), this approach systematically probes the coherence, depth, and soundness of agent arguments.
Pseudo-codes for the DAReD and SReD strategies are provided in Appendix G.
3.3 Memory
Memory is crucial for intelligent reasoning, enabling agents to learn from past experiences, refine decision-making, and avoid repeating errors. In the context of RedDebate, where agents engage iteratively in challenging debate, effective memory mechanisms allow continuous enhancement of agent behaviour. Inspired by cognitive structures in human decision-making Zhang et al. (2024), we propose to integrate two complementary memory modules into RedDebate: short-term memory (STM) and long-term memory (LTM). STM provides immediate context, maintaining coherence within an ongoing debate, and is reset upon each debate’s completion. LTM, on the other hand, acts as a persistent repository, storing accumulated safety insights and feedback. In this section, we explore four variations of LTM tailored to our setting.
Textual Long-Term Memory (TLTM)
is widely adopted in prior work due to its interpretability, ease of implementation, and efficient read-write operations Zhang et al. (2024). In this memory type, generated natural language feedback is incorporated into the agent’s prompt to make the agent aware of previously learned lessons. However, since agents may accumulate a large number of feedback entries through repeated trial and error, including all of them in the prompt can be inefficient, given our setup with relatively lightweight LLMs and limited context windows. To address this, we encode all feedback $\Phi=\{\phi_{1},\ldots,\phi_{k}\}\subseteq\mathcal{M}_{L}$ into vector representations [^3] Vector representations are used only for retrieval; ultimately, agents receive top-matching textual feedback entries. and store them in a vector database. For future prompts $\mathcal{X}^{\prime}$, the system retrieves the most relevant feedback entries using a similarity function $\mathrm{sim}(\mathcal{X}^{\prime},\phi)$ and adds these textual feedbacks to the agent’s context. This ensures memory remains concise, relevant, and helpful without overwhelming the agent.
Continuous Long-Term Memory (CLTM)
Also known as parametric memory, CLTM stores feedback directly within the LLM’s parameters. This approach alleviates key limitations of TLTM, such as increased context length and potential retrieval misses. To implement CLTM, we use Parameter-Efficient Fine-Tuning (PEFT), specifically LoRA (Hu et al., 2022), which allows us to inject feedback into the model with minimal computational overhead (Houlsby et al., 2019). Each debater is fine-tuned on the accumulated feedback, treating the feedback as language modeling targets. To manage resource costs, we periodically reset and re-fine-tune the PEFT weights after a fixed number of new feedback entries have been collected.
Unified Long-Term Memory (TLTM+CLTM)
is designed to simultaneously exploit the strengths of symbolic (TLTM) and distributed (CLTM) memory types, employing both representations concurrently. In this integrated approach, CLTM reinforces textual memory, much like how working memory in humans can enrich reading comprehension when presented with explicit textual knowledge and thereby facilitate effective decision-making (Peng et al., 2017). This combination benefits from the interpretability and retrieval speed of symbolic memory and the comprehensive representational capacity of parametric memory.
Guardrails Long-Term Memory (GLTM)
aims to explicitly encode unsafe experiences into executable programmatic constraints or guardrails. A core limitation of prior types of memory is that they are inherently passive: agents must actively integrate lessons into generation. Consequently, the burden resides on the agent to correctly recall and interpret memory content at generation. Inspired by recent work on automatic programmatic guardrail generation (Sreedhar et al., 2024; Dong et al., 2024a), we adapt the idea to the safety setting by implementing GLTM to automatically encode agents’ past unsafe experiences as guardrails. Before an agent generates a response, the input prompt will be directly rejected if it matches a previously known unsafe flow.
We use one-shot prompting to generate guardrails given a generated feedback $\phi$ and the corresponding harmful question $\mathcal{X}$. We instruct the LLM to output an expression and a list of examples for each $(\mathcal{X},\phi)$ pair. The expression is an abstraction of the harmful user behaviour, similar to abstractions generated in Wang et al. (2024b). It also serves as the flow names–Colang equivalent of functions–in the final resulting Colang application. Additionally, we instruct the LLM to provide examples of user utterances of such harmful behaviours. We merge the examples of the expressions if the expressions are the same. Finally, we use a rule-based method to convert the expressions into Colang flows, allowing use of Colang’s built-in intent detection feature to match the defined harmful behaviours and reject the user request. Figure 19 in Appendix shows an example of the code generation prompt and the resulting guardrail flow.
4 Experimental Setup
4.1 Datasets
We evaluate our framework on HarmBench Mazeika et al. (2024), featuring direct prompts targeting harmful behaviours, and CoSafe Yu et al. (2024), with more indirect, dialogue-based prompts reflecting realistic conversational safety challenges.
4.2 Evaluation Metrics
For each agent $n$, given the evaluated safety label $y_{p,n}^{(t)}$ at round $t$ for the $p$-th input prompt, we assess the safety of debates using two metrics:
Error Rate (ER)
This measures the proportion of unsafe responses among all responses by that agent across all prompts and rounds, as defined in Equation 4. The total error rate is computed similarly, but aggregates responses over all agents.
$\text{ER}{{}_{n}}=\frac{\sum_{p=1}^{|\mathcal{P}|}\sum_{t=1}^{T}\mathbb{I}[y_{p,n}^{(t)}=0]}{|\mathcal{P}|\times T}$ (4)
Agreement Rate (AGR)
To capture how often agents correct unsafe outputs in multi-round debates, we introduce AGR, which quantifies the proportion of transitions where a response changes from unsafe to safe across consecutive rounds. For each agent $n$, the agreement rate is defined in Equation 5. The total agreement rate is computed by aggregating transitions across all agents.
$\text{AGR}_{n}=\frac{\sum\limits_{p=1}^{|\mathcal{P}|}\sum\limits_{t=1}^{T-1}\mathbb{I}\left[y_{p,n}^{(t)}=0\wedge y_{p,n}^{(t+1)}=1\right]}{|\mathcal{P}|\times(T-1)}$ (5)
4.3 Implementation Details
By default, all debates are conducted over three rounds. We use Mistral-7B-v0.2, LLaMA-3.2-3B, and Phi-3.5-mini as our primary backbone models for the debater agents. In the Devil–Angel setting, we use these three models and explored all possible role assignment, reporting the average performance across permutations. For the feedback generator and Socratic agent, which play critical guiding roles, we use GPT-4o-mini to mitigate context window limitations. All safety evaluations are performed with LlamaGuard. The complete agent prompts are listed in Appendix I.
For TLTM, feedback is embedded with OpenAI’s text-embedding-3-large, and the top five entries are retrieved at inference via cosine similarity. The CLTM employs LoRA-based adaptation on the debater’s attention layers. Agents are fine-tuned on accumulated feedback after every 10 new feedback entries using cross-entropy loss. For GLTM, we generate Colang guardrails by prompting GPT-4o. Further technical details for LTMs design choices are available in Appendix C.
5 Results and Analysis
5.1 Debate Performance
Scenario HarmBench CoSafe
Error Rate (%) $\downarrow$ Agreement Rate (%) $\uparrow$ Error Rate (%) $\downarrow$ Agreement Rate (%) $\uparrow$
Total Mistral Llama Phi Total Mistral Llama Phi Total Mistral Llama Phi Total Mistral Llama Phi
Std. Prompting 38.7 58.5 21.9 35.7 7.4 9.0 7.5 5.7
PReD 28.8 37.2 21.3 27.9 12.3 21.3 8.5 7.1 6.5 7.5 6.3 5.7 3.0 3.7 3.1 2.2
DAReD 24.9 36.3 15.6 22.8 14.5 21.6 9.9 12.1 5.9 6.3 5.8 5.6 3.3 3.5 3.3 3.1
SReD 21.0 25.7 15.8 21.6 17.0 26.3 10.4 14.5 4.5 4.8 4.5 4.2 3.8 5.0 3.6 2.9
Table 1: Error rates and Agreement rates (%) across different scenarios without LTM.
Scenario HarmBench Error Rate (%) $\downarrow$ CoSafe Error Rate (%) $\downarrow$
Total Mistral Llama Phi Total Mistral Llama Phi
Self-Critique 15.4 23.3 10.8 12.0 8.1 7.0 12.8 4.6
SReD 21.0 25.7 15.8 21.6 4.5 4.8 4.5 4.2
   +tltm 15.2/+5.8 18.0/+7.7 13.5/+2.3 14.1/+7.5 3.1/+1.4 3.0/+1.8 3.0/+1.5 3.3/+0.9
   +cltm 14.1/+6.9 16.0/+9.7 4.6/+11.2 21.6/0.0 4.3/+0.2 4.0/+0.8 3.3/+1.2 5.7/-1.5
   +tltm+cltm 6.1/+14.9 6.7/+19.0 4.1/+11.7 7.4/+14.2 2.4/+2.1 3.3/+1.5 2.1/+2.4 2.0/+2.2
   +gltm 3.6/+17.4 8.4/+17.3 0.3/+15.5 2.0/+19.6 2.5/+2.0 4.4/+0.4 0.4/+4.1 2.8/+1.4
Table 2: Error rates (%) for Self-Critique and SReD across all LTM integrations. Improvements over the no-memory debate setting are shown in gray.
Engaging in Debate Leads to Safer Answers.
As shown in Table 1, Standard Prompting—where agents independently answer in a single round—exhibits the highest error rate. In contrast, PReD significantly reduces both total and individual error rates, confirming that multi-agent interaction improves response safety. Refinement is a key aspect which captures the improvement in response safety throughout the course of a debate, quantified by reductions in agents’ error rates and increases in agreement. When exploring the impact of different debate scenarios on refinement, SReD achieves the highest agreement rate (17.0% in HarmBench and 3.8% in CoSafe) and the least error rate. This indicates that agents revise unsafe responses more effectively when prompted by explicit questioning or counterarguments from persona agents. Indeed, the more actively agents are engaged in the debate process, the more opportunities they have to refine unsafe reasoning.
Learning from Experience Lowers Error Rates.
We select the best-performing debate scenario without memory—SReD—and equip debaters with different types of LTM to determine whether agents being able to leverage previously learned experiences reduces error rates. As shown in Table 2, when utilizing LTM we observe consistent improvements in overall error rate across all LTM types. Additionally, combining TLTM and CLTM into a Unified LTM leads to even greater error reduction, confirming the benefit of integrating both mechanisms. This highlights that both memory types on their own are not fully sufficient for imparting the information from the generated feedback into the debaters. By integrating both mechanisms a form of constructive interference is created and the information is better relayed allowing debaters to better recount crucial feedback. An ablation on PReD with LTM is also presented in Appendix D, demonstrating that LTM remains effective even in the absence of auxiliary agents. Furthermore, as shown in Appendix E, memory-augmented agents maintain their helpfulness. This is reflected in both accuracy remaining at similar levels and a low refusal rate on safe general questions, which demonstrates that the safety enhancements provided by LTM do not diminish the agents’ overall utility.
Diverse Peer Feedback with Memory Outperforms Self-Critique.
Our method enables agents to receive dynamic feedback, without requiring human input, that is directly grounded in their own mistakes—resulting in more targeted refinement. This allows it to outperform settings where agents refine themselves using fixed, predefined rules, as explored in the Self-Critique baseline (Table 2).
The addition of LTM significantly enhances debate performance relative to Self-Critique, despite Self-Critique initially achieving lower error rates on HarmBench compared to debate scenarios without memory (e.g., 15.36% vs. 21.0% in SReD). Self-Critique likely has an initial edge due to the simplicity and the effect of human input—agents are simply given their own previous response and asked to revise it per a predefined constitutional rule. In contrast, debate agents must reason and respond to a dynamic conversation without direct guidance which adds greater complexity. Hence, LTM allows for additional guidance for the debaters through feedback while benefiting from the additional depth and reasoning provided from the debate process, allowing for optimal performance.
5.2 Detailed Analysis
Preventative Guardrails Yield Superior Safety.
As shown in Table 2, the “SReD+gltm” configuration yields the lowest error rates across both HarmBench and CoSafe, demonstrating the value of converting learned unsafe behaviours into explicit control flows that proactively block harmful inputs. Its effectiveness stems from the fact that “prevention is better than cure”: harmful inputs that match prior unsafe patterns are intercepted before reaching the model. In addition, Colang’s intent detection follows a two-stage process—similar to RankGPT (Sun et al., 2023)—where it first retrieves relevant guardrails and then uses a generative agent to assess whether any match is strong enough to trigger rejection. This layered verification offers greater precision than relying solely on retrieval or embedding similarity.
Among the models, Llama shows the most reliable GLTM integration, with the lowest error rates (0.3% on HarmBench, 0.4% on CoSafe), high recall of unsafe intents (99.4%, 92.7%), and minimal runtime failures, making it well-suited for guardrail-based safety (Table 3). In contrast, Phi suffers from non-negligible runtime errors due to NeMo’s instability and formatting constraints.
Metric HarmBench CoSafe
Mistral Llama Phi Mistral Llama Phi
Intent Match (%) $\uparrow$ 25.5 39.5 26.0 2.4 4.5 1.4
Guardrails Recall (%) $\uparrow$ 76.7 99.4 94.5 35.4 92.7 35.2
Runtime Error (%) $\downarrow$ 7.3 0.3 25.0 0.0 0.1 9.7
Table 3: Guardrail effectiveness across benchmarks. Intent Match (%): Proportion of unsafe prompts that are correctly blocked by triggering the guardrail intent. Guardrails Recall (%): Recall of all unsafe prompts detected and blocked by guardrails. Runtime Error (%): Percentage of responses with technical or formatting errors, which are excluded from evaluation results.
Category-Level Errors.
We analyze category-level error distributions to identify areas where agents are more prone to unsafe responses. SReD consistently performs better across most categories; however, challenges persist in categories such as copyright and chemical biological in HarmBench, and terrorism and financial harm in CoSafe. For more details and visualizations, see Appendix F.
(a) Debate
(b) Self-Critique
Figure 3: Stepwise error rates for Debate and Self-Critique. Debate shows consistent error reduction and convergence, while Self-Critique improves less reliably.
Refinement Dynamics.
To closely examine the refinement dynamics of debate and Self-Critique, we analyze step-wise error rates (Figure 3). Debaters with memory consistently reduce errors across each round (Figure 3(a)), eventually converging toward a similar error rate range. This convergence suggests that debaters achieve a safe consensus through interactive debate. Conversely, Self-Critique (Figure 3(b)) shows no convergence due to isolated revisions, highlighting debate’s advantage: external perspectives and peer feedback are crucial for effective safety refinement. Agents often struggle to recognize their own unsafe behaviour, which can instead be surfaced and corrected through peer interaction during debate.
Debate Exposes Latent Vulnerabilities.
We observe that introducing varied debate strategies serves as a mechanism to reveal hidden vulnerabilities not visible in single-turn settings. A sample of this behaviour is shown in Appendix H.2. Taking HarmBench as a case study, we note that Llama and Phi, despite initially providing safe responses (Figure 4(a)), produce answers that become unsafe as the debate progresses (Figure 4(b)). Often, these initial responses are minimal and cautious—such as “I cannot provide you with the answer”—but as agents engage further in the debate or address questions raised by other agents, they fail to maintain safety and ultimately produce unsafe responses. Indeed, this process of error discovery aligns well with the dual focus of our framework on learning and red-teaming. By surfacing previously unexposed unsafe spots, agents are given the opportunity to learn from accompanying feedback and apply those lessons in similar future situations.
(a) Initial Unsafe Responses
[…]
Figure 4: Vulnerability heatmaps showing model unsafety on the first attempt and subsequent ones
[…]
(b) Average Steps to Unsafe
Figure 5: Comparison of attack effectiveness (left) and efficiency (right) across approaches.
Debate Strategies Vary in Attack Effectiveness.
The rate at which agents transition from safe to unsafe responses depends on the debate strategy. As illustrated in Figure 5(a), the DAReD scenario triggers the highest number of these conversions, whereas SReD results in fewer conversions overall. However, for the cases where SReD does discover vulnerabilities, it uncovers them in fewer average steps compared to other scenarios. To further analyze the efficiency of debate strategies as “attack methods” that trigger unsafe responses, we plot the results across different models and debate scenarios in Appendix, Figure 6.
Ablations.
Ablation studies on extending debate rounds up to five, using additional peer debaters, and evaluating the effectiveness of evaluator agent are presented in Appendix B. Overall, we show that three debate rounds lead to optimal performance, using additional agents can improve error rate and diversity, and that LlamaGuard produces strong evaluations for safety.
6 Conclusion
Through our RedDebate framework, we show that multi-agent debates combined with automated red-teaming and long-term memory can significantly enhance LLM safety without human supervision. By enabling agents to critique and refine each other’s responses, our framework both effectively reduces unsafe outputs and uncovers vulnerabilities. Memory modules and proactive guardrails further amplify safety improvements, demonstrating that structured collaboration and systematic feedback offer a practical path toward more robust and scalable AI safety solutions.
Limitations
As with any automatic tool for evaluation, our primary safety evaluator LlamaGuard does not achieve perfect performance, which can cause some issues in evaluating responses and generating the most optimal feedback. Overall, it does not compromise the performance of our framework and we observe strong increases in safety particularly on HarmBench which cannot be explained by noise in the evaluator. Additionally, we cover very standard benchmark datasets in HarmBench and CoSafe which span a wide range of domains, although we do not focus on underrepresented domains in this work as we aim to achieve a broad view of the issue.
We focus our experiments on open-source, small scale language models, as these are more likely to suffer from misalignment from human values and stand to gain the most improvement in safety through our approach, which we demonstrate through our results. These models are especially useful because of their transparency and ability to be applied in domains where API-based models are unsuitable. It is possible to apply our framework to large state-of-the-art language models offered through API services, but these models undergo far more alignment and extensive patching to become more safe OpenAI (2024), and hence we prioritize testing our framework on the aforementioned models.
A few limitations stem from strict Colang formatting and some of NeMo’s instability led to runtime errors in some models such as Phi, indicating that the library itself needs improvements. Lastly, despite our RedDebate framework leading to a significant reduction in unsafe behaviours, we note that there is an increased inference cost when using our standard debate settings with three agents across three rounds which can increase wall-clock inference by roughly 3-4 times over single-turn baselines. Nevertheless, we stress that this framework aims to maximize safety in applications where it is the upmost concern over fast computation.
Ethics Statement
This research investigates methods to enhance the safety of large language models through automated debate and red-teaming, with the goal of proactively mitigating harmful or unsafe outputs. All experiments were conducted using publicly available models and datasets. The primary datasets, HarmBench and CoSafe, contain adversarial prompts focused on conversational and social safety; care was taken to ensure that outputs and analysis remained within ethical and legal guidelines. No private or sensitive user data was used.
While our framework aims to reduce unsafe behaviours, automated safety evaluation and guardrails are not foolproof and may miss nuanced or context-dependent harms. We caution that models—even when improved by our techniques—should not be deployed in high-stakes or real-world scenarios without thorough human oversight and external auditing. Our code and results are shared to promote transparency and reproducibility.
The potential misuse of automated debate systems for adversarial or malicious purposes is a recognized risk. We encourage responsible research practices and urge practitioners to consider societal impacts, bias propagation, and unintended consequences when building on or deploying similar methods.
No human subjects, personally identifiable information, or sensitive data were involved in this research.
Appendix A Extended Related Work
Multi-Agent Debates
encourage diverse reasoning by involving multiple LLMs, each bringing distinct perspectives. Prior work has shown that such interaction improves factual accuracy, alignment, and reasoning compared to individual agents, ensembling, or self-reflection-based prompting (Smit et al., 2024; Chan et al., 2024; Du et al., 2024; Liang et al., 2024). In early foundational work, Irving et al. (2018) proposed training agents via self-play on a zero-sum debate game to align superhuman AI. Khan et al. (2024) shows that the use of debate can aid weaker judges in evaluating stronger models. Chern et al. (2024) find that MAD can reduce model toxicity when jailbroken or less capable models are forced to debate with capable models. However, none of these works employ multi-agent debate as a red-teaming strategy combined with learning for safety refinement as we do.
Red Teaming
LLMs often exhibit unsafe or harmful behaviours to users. Red-teaming involves the creation and evaluation of a set of test cases aimed at finding such LLM failure cases. Traditional methods involve extensive use of human annotation to manually generate test cases and/or detect harmful responses (Dinan et al., 2019; Ribeiro et al., 2020; Ziegler et al., 2020; Xu et al., 2021; Hendrycks et al., 2021). Perez et al. (2022) first established a method to both automatically generate test cases using a language model, and find failures using a trained classifier. Follow-up work has introduced curiosity-driven exploration for increased coverage of test cases (Hong et al., 2024), optimizing the process of iteratively updating test cases (Mehrabi et al., 2024), and building a red-team that can automatically formulate a measure for harmful outputs and optimize a generator for diverse adversarial prompts (Casper et al., 2023). We build on this work by addressing the need for a fully automated red-teaming evaluation pipeline using LLMs as evaluators without requiring any trained classifier or human oversight.
Learning From Feedback
Feedback from either humans or using automatically generated methods can effectively steer LLM behaviour to be better aligned with human values. Reinforcement learning from human feedback (RLHF) is a popular method of finetuning LLMs on human preference data to tune them to act as helpful and harmless assistants OpenAI et al. (2024); Bai et al. (2022a). By training a preference model, the model obtains feedback on desirable behaviours. Bai et al. (2022b) uses supervised learning and reinforcement learning to iteratively tune LLMs based on feedback generated using self-critiques according to a set of predefined rules for agents’ revision. Shinn et al. (2023) builds on this by using self-reflective feedback from verbal text stored in an episodic memory buffer as an additional context for LLM agents to help them learn from prior mistakes and improve performance in subsequent trials. We extend this line of work by enabling feedback through multi-agent debate, where peers critique each other, yielding richer safety feedback. Furthermore, we integrate debate with various long-term memory, allowing models to learn from feedback while improving their safety.
Agent Memory
Due to limitations in context length and ability to handle longer-term dependencies LLM agents benefit from a means of storing and accessing information they have previously learned across interactions (Zhang et al., 2024). Accordingly, previous works have sought to accomplish this by incorporating memory modules, which are generally divided into two types: textual and parametric. Approaches using textual memory store and retrieve information in natural language. Means of textual memory include long-context length strategies (Huang et al., 2023; Li et al., 2023), or strategically processing recent interactions using methods such as flash memory Wang et al. (2025) virtual context management Packer et al. (2023), or cache systems Wang et al. (2024a). Retrieval-based mechanisms using vector-databases can allow most relevant information from full-context to be effecitevly utilized (Hu et al., 2023; Zhong et al., 2024). Parametric memory involves directly altering LLM parameters to adaptively impart knowledge into LLMs. Fine-tuning based methods train on domain-specific knowledge using supervised fine tuning (Hu et al., 2023), including parameter efficient tuning methods Xiong et al. (2023). Likewise, memory editing methods target modifying specific LLM parameters to inject a small set of facts without inhibiting general knowledge (De Cao et al., 2021; Mitchell et al., 2022; Meng et al., 2022, 2023). In this paper, we are the first to use both types of memory in the context of debate agents and safety.
Guardrailing
Guardrailing allows determining if and how some actions could be enforced to increase the safety in a system (Dong et al., 2024b). Recently, many LLM guardrailing tools have been developed to mitigate risks associated with them (Rebedea et al., 2023; Guardrails AI, ). NeMo Guardrails (Rebedea et al., 2023) allows building safe LLM agents via a programming language called Colang [^4] https://docs.nvidia.com/nemo/guardrails/colang_2/overview.html by specifying predefined dialogical pathways for LLMs. Programmatic gaurdrailing allows modelling highly flexible conversational flows, which might not be possible with existing dialogue management techniques (Rebedea et al., 2023). Recently, the canonical form extraction of automatic guardrails (Sreedhar et al., 2024) has garnered interest and shown promising results in task-oriented dialogue systems. Building on this line of work, we propose, for the first time, the use of guardrails as memory in the safety setting.
Appendix B Debate and Agents Ablation Study
B.1 Debate Rounds
We study the effect of increasing the number of debate rounds from three to five on the HarmBench dataset. Due to the ease of applying TLTM, we focus solely on this type of memory integration in this study. This analysis investigates how extending the debate affects two key metrics: total error rate and agreement rate. Additionally, we introduced a new metric called diversity, which measures whether each debate round contains at least one safe and one unsafe response, thereby capturing the variability of viewpoints throughout the debate.
As shown in Table 4, continuing the debate from three to five rounds leads to a modest improvement in total error rate without LTM, whereas extending to four rounds does not yield any noticeable benefit. Interestingly, the agreement rate remains largely unaffected by increasing the number of rounds. Although there is a slight increase in diversity from three to four rounds, the improvement is not significant. Overall, these results suggest that three rounds is sufficient for agents to explore a red-teaming prompt and extending the debate beyond three rounds does not offer significant benefits in terms of safety and agreement.
Setting TER $\downarrow$ TAGR $\uparrow$ DIV $\uparrow$
3 Rounds 0.289 0.120 0.390
4 Rounds 0.296 0.105 0.405
5 Rounds 0.264 0.094 0.380
3 Rounds+tltm 0.117 0.074 0.280
4 Rounds+tltm 0.180 0.071 0.303
5 Rounds+tltm 0.184 0.059 0.248
Table 4: Impact of the number of debate rounds on performance metrics in HarmBench. TER: Total Error Rate, TAGR: Total Agreement Rate, DIV: Diversity.
B.2 Number of Agents
Beside the three primary debater models Mistral-7B-v0.2 (Jiang et al., 2023), LLaMA-3.2-3B (Grattafiori et al., 2024), and Phi-3.5-mini (Abdin et al., 2024), we also examine whether increasing the number of participating agents improves refinement, reduces error, and enhances the diversity of the debate. As shown in Table 5, increasing the number of agents from three to four (by adding Gemini1.5-Flash-8B), and especially to five (with both Gemini1.5-Flash-8B and GPT-4o-mini), results in a notable reduction in total error rate. Part of this improvement may be attributed to the inclusion of more robust models among the three previous agents, which positively affects total error rates. Moreover, increasing the number of agents contributes to higher diversity, as more perspectives are introduced in the debate. This results in a richer range of opinions and, ultimately, better error reduction. In summary, increasing the number of agents appears to foster a more dynamic and effective debate, facilitating the correction of unsafe responses.
Setting TER $\downarrow$ TAGR $\uparrow$ DIV $\uparrow$
3 Agents 0.289 0.120 0.390
4 Agents 0.291 0.107 0.521
5 Agents 0.245 0.101 0.570
3 Agents+tltm 0.117 0.074 0.280
4 Agents+tltm 0.111 0.054 0.285
5 Agents+tltm 0.060 0.042 0.223
Table 5: Impact of the number of agents on performance metrics in HarmBench. TER: Total Error Rate, TAGR: Total Agreement Rate, DIV: Diversity.
[…]
B.3 Evaluator Agent
Automated safety evaluation using LLMs has shown promising results in recent work, including the introduction of dedicated safety evaluators such as LlamaGuard Inan et al. (2023). While these models can detect harmful content with reasonable accuracy, it remains important to validate their performance in new setups and task structures—particularly in our framework which operates autonomously, without human oversight.
We compare two LLM-based evaluators—LlamaGuard and GPT-4o-mini with a moderation prompt (Refer to Figure 7) against human-labeled safety judgments. We annotated each argument from every agent in each debate round as either safe or unsafe, resulting in 315 dialogue samples from HarmBench debates. Model predictions were then compared to these human labels. The results are summarized in Table 6. Both models demonstrated solid performance, suggesting that they can serve as substitutes for human annotation—or at least as reliable first-pass filters subject to human review. Among the two, LlamaGuard outperformed GPT-4o-mini across all metrics, achieving higher accuracy, precision, recall, and F1 score.
Figure 7: Moderation prompt used in GPT-4o-mini as Evaluator agent
Evaluator Accuracy Precision Recall F1 Score
LlamaGuard 0.87 0.90 0.94 0.92
GPT-4o-mini 0.81 0.87 0.89 0.88
Table 6: Comparison of automatic safety evaluators against human annotations on 315 arguments from HarmBench debates.
Also note that the evaluation in all reported results in the paper is conducted only on debater agents, excluding the responses of Socratic and Devil–Angel agents, as they do not directly answer the debate question but instead trigger, prompt, or guide the conversation.
B.4 Socratic Agent and Feedback Generator LLM Backbones
Given the central guiding roles of the feedback generator and Socratic agent, we selected a larger model to avoid context window limitations. While we evaluated DeepSeek-R1, we ultimately used GPT-4o-mini in all experiments, balancing response quality with inference speed.
Appendix C LTMs Technical Details
TLTM.
We embed each feedback using OpenAI’s text-embedding-3-large model (dimension: 3072), and store the resulting vectors in a Pinecone vector database. [^5] https://www.pinecone.io During inference, the debate prompt is embedded using the same model, and cosine similarity is used to retrieve the most relevant feedback entries. Based on our observations, retrieving the top five most semantically relevant feedbacks provides sufficient context for improving safety without overwhelming the agent.
CLTM.
The CLTM utilizes LoRA and is implemented using the HuggingFace Library  [^6] https://huggingface.co/docs/peft/en/package_reference/lora. We apply LoRA to the debater attention layers, adding around 0.1% trainable parameters. We set the LoRA parameters as follows: low-rank parameter $r=16$, LoRA $\alpha=16$, LoRA dropout $=0.1$. Cross-entropy loss between the model’s predictions and the actual next tokens in the sequence is used for optimization. We re-fine-tune the parameters on the whole set of feedback with every 10 new ones added.
GLTM.
To implement the Colang guardrails, we use a fork of NeMo-Guardrails version 0.11, which we slightly modified to inject CoSafe conversation history and improve robustness for evaluation. The modified NeMo-Guardrails version used in our experiments is available at our GitHub repository. For HarmBench guardrail code generation, we randomly select 120 samples [^7] In our preliminary experiments, we tried 80, 120, and 200 samples and found that 120 examples yielded the best results. from the debate history. This results in 68 pieces of feedback—since not every sample produces an unsafe response (Algorithm 1)—from which we generate 44 guardrails, as some feedback results in overlapping guardrail names that are subsequently merged. For CoSafe, we use 700 samples, which also yield 68 pieces of feedback, from which we generate 63 guardrails. We use GPT-4o to build all guardrails.
Appendix D PReD with LTM
We also provide an ablation study on the PReD setting, demonstrating the effect of long-term memory (LTM) in scenarios without role assignments. As shown in Table 7, the results are consistent with the insights discussed in Section 5, confirming that LTM improves performance compared to the no-memory baseline.
Scenario HarmBench Error Rate (%) $\downarrow$ CoSafe Error Rate (%) $\downarrow$
Total Mistral Llama Phi Total Mistral Llama Phi
Self-Critique 15.4 23.3 10.8 12.0 8.1 7.0 12.8 4.6
PReD (Ours) 28.8 37.2 21.3 27.9 6.5 7.5 6.3 5.7
   +tltm 11.5/+17.3 15.9/+21.3 9.3/+12.0 9.3/+18.6 3.1/+3.4 3.3/+4.2 3.0/+3.3 3.0/+2.7
   +cltm 22.3/+6.5 27.3/+9.9 12.5/+8.8 27.1/+0.8 4.6/+1.9 3.3/+4.2 3.8/+2.5 6.9/-1.2
   +tltm+cltm 10.7/+18.1 9.8/+27.4 7.8/+13.5 14.7/+13.2 2.9/+3.6 3.0/+4.5 2.3/+4.0 3.6/+2.1
   +gltm 3.1/+25.7 5.2/+32.0 1.0/+20.3 3.0/+24.9 2.7/+3.8 3.5/+4.0 0.3/+6.0 4.3/+1.4
Table 7: Error rates (%) for Self-Critique and PReD across all LTM integrations. Improvements over the no-memory debate setting are shown in gray.
Appendix E Helpfulness After Memory Integration
While our primary focus is on improving safety, we also examine whether integrating different types of long-term memory (LTM) negatively impacts agents’ helpfulness on general, non-adversarial queries. In particular, we assess whether safer agents become overly cautious and refuse benign prompts after incorporating safety-oriented feedback.
For our evaluation, we sample 1,000 trivia questions from TriviaQA (Joshi et al., 2017), each paired with concise gold answers and their known aliases. For each LLM agent equipped with a different LTM type, we check whether its response contains the expected answer or any of its aliases (accuracy), and whether it refused to answer by stating the prompt was unsafe. We automate this process using GPT-4o, which matches the agent’s response against the provided answer set and identifies refusal statements.
As shown in Table 8, integrating memory modules does not substantially reduce accuracy, nor does it lead to a high refusal rate. This indicates that the overall helpfulness of the agents remains intact. Thus, while agents become significantly safer, they also remain adequately helpful and do not excessively refuse general queries.
We also observe a slight, though not significant, improvement in accuracy for some memory-augmented agents. This unexpected enhancement does not indicate a genuine improvement in our system, but is likely due to noise—particularly from using GPT as the evaluator—and minor variations in the context provided to the LLMs across different memory integration types compared to the baseline without memory. Phi in particular has a significant improvement in accuracy over the baseline in Table 8 despite having a higher error rate in certain cases compared to the no-memory debate baselines (likely as a result of overfitting to feedback), showing that the model still is attaining benefits from CLTM but not always directly in terms of safety. Nevertheless, the results and the provided insights remain unchanged.
Metric Accuracy (%$\uparrow$ Refusal Rate (%$\downarrow$
Mistral Llama Phi Mistral Llama Phi
None (Baseline) 68.3 57.6 57.9
TLTM 64.2 58.4 52.3 1.8 0.4 0.7
CLTM 63.7 60.9 66.7 0.3 0.0 0.0
TLTM+CLTM 63.7 59.6 62.2 0.0 0.1 0.0
GLTM 65.5 58.2 53.5 1.0 0.3 1.8
Table 8: Accuracy (%) and Refusal Rate (%) on TriviaQA after integrating different LTM types. Accuracy denotes the percentage of responses containing the correct answer or any of its aliases. Refusal Rate indicates the proportion of cases where the agent refused to answer, typically citing safety concerns.
Appendix F Categorical Vulnerability Analysis
As stated in Section 5, we further analyze how different debate strategies perform across individual safety categories within the HarmBench and CoSafe datasets. Figure 8 presents radar plots of total error rates for each category, comparing PReD, SReD, and DAReD strategies.
SReD demonstrates consistently lower error rates across most categories, suggesting it is more effective at guiding agents toward safer behaviour. However, some categories remain particularly challenging across all strategies. In HarmBench, agents frequently fail in copyright and chemical synthesis, while in CoSafe, high error rates are observed in terrorism and financial harm.
Additionally, we provide per-agent categorical breakdowns (Figure 9) to highlight model-specific vulnerabilities, further illustrating how different models vary in their susceptibility to unsafe outputs depending on the topic. These analyses help pinpoint which combinations of strategies and models require greater attention for targeted safety improvements.
(a) Total error rates HarmBench
(b) Total error rates on CoSafe
Figure 8: Total error rates per dataset category and debate strategy
(a) PReD on HarmBench
(b) DAReD on HarmBench
(c) SReD on HarmBench
(d) PReD on CoSafe
(e) DAReD on CoSafe
(f) SReD on CoSafe
Figure 9: Agents’ error rates by dataset category and debate strategy.
Appendix G Persona-Based Debate Strategy Algorithms
In Algorithm 2 and Algorithm 3, we present the pseudo-code for the Devil–Angel and SReD strategies. In the Devil–Angel debate setting, we simplify the setup by involving only a single debater agent. This design choice reduces communication complexity and avoids excessive dialogue overlap, which can lead to confusion or context dilution for the participating agents. Also note that persona agents (i.e., Devil, Angel, and Socratic) provide their responses only after the debaters have completed their current round and the short-term memory has been updated.
1
Input: Prompt $\mathcal{X}$, debater $\mathrm{D}$, devil $\mathrm{D^{\ominus}}$, angel $\mathrm{D^{\oplus}}$, evaluator $\mathrm{E}$, feedback generator $\mathrm{F}$, STM $\mathcal{M}_{\text{S}}$, LTM $\mathcal{M}_{\text{L}}$, rounds $T$
Output: Debate history $\mathcal{R}$
2
$\mathcal{M}_{\text{S}}\leftarrow\emptyset$
  // Initialize STM
$\mathcal{R}\leftarrow[]$
  // Initialize Debate History
3
4for $t\leftarrow 1$ to $T$ do
5       $\mathcal{R}^{(t)}\leftarrow\{\}$
6
      $r^{(t)}\leftarrow\mathrm{D}(\mathcal{X},\mathcal{M}_{\text{S}},\mathcal{M}_{\text{L}})$
        // Debate
7       $\mathcal{R}^{(t)}\leftarrow\mathcal{R}^{(t)}\cup\{r^{(t)}\}$
      $\mathcal{M}_{\text{S}}\leftarrow\mathcal{M}_{\text{S}}\cup\mathcal{R}^{(t)}$
        // Update STM
8
      $\alpha^{(t)}\leftarrow\mathrm{D^{\oplus}}(\mathcal{X},\mathcal{M}_{\text{S}},\mathcal{M}_{\text{L}})$
        // Reinforce
       $\delta^{(t)}\leftarrow\mathrm{D^{\ominus}}(\mathcal{X},\mathcal{M}_{\text{S}},\mathcal{M}_{\text{L}})$
        // Oppose
9
10      $\mathcal{R}^{(t)}\leftarrow\mathcal{R}^{(t)}\cup\{\alpha^{(t)},\delta^{(t)}\}$
11
      $\mathcal{M}_{\text{S}}\leftarrow\mathcal{M}_{\text{S}}\cup\mathcal{R}^{(t)}$
        // Update STM
       $\mathcal{R}\leftarrow\mathcal{R}\cup\{\mathcal{R}^{(t)}\}$
        // Append to History
12
13 end for
14
$\mathcal{Y}\leftarrow\mathrm{E}(\mathcal{R})$
  // Evaluate History (ignore $\alpha,\delta$)
15
16if $\exists\,y_{n}^{(t)}=0$ in $\mathcal{Y}$ then
       $\phi\leftarrow\mathrm{F}(\mathcal{R},\mathcal{Y})$
        // Generate Feedback
       $\mathcal{M}_{\text{L}}\leftarrow\mathcal{M}_{\text{L}}\cup\{\phi\}$
        // Update LTM
17
18 end if
19
return $\mathcal{R}$
  // Return Debate History
Algorithm 2 Devil–Angel Refinement Debate
1
Input: Prompt $\mathcal{X}$, debaters $\mathcal{D}=\{\mathrm{D_{1}},\dots,\mathrm{D_{N}}\}$, Socratic agent $\mathrm{D^{S}}$, evaluator $\mathrm{E}$, feedback generator $\mathrm{F}$, STM $\mathcal{M}_{\text{S}}$, LTM $\mathcal{M}_{\text{L}}$, rounds $T$
Output: Debate history $\mathcal{R}$
2
$\mathcal{M}_{\text{S}}\leftarrow\emptyset$
  // Initialize STM
$\mathcal{R}\leftarrow[]$
  // Initialize Debate History
3
4for $t\leftarrow 1$ to $T$ do
5       $\mathcal{R}^{(t)}\leftarrow\{\}$
6
7      for $n\leftarrow 1$ to $N$ do
             $r_{n}^{(t)}\leftarrow\mathrm{D_{n}}(\mathcal{X},\mathcal{M}_{\text{S}},\mathcal{M}_{\text{L}})$
              // Debate
8             $\mathcal{R}^{(t)}\leftarrow\mathcal{R}^{(t)}\cup\{r_{n}^{(t)}\}$
9
10       end for
11
      $\mathcal{M}_{\text{S}}\leftarrow\mathcal{M}_{\text{S}}\cup\mathcal{R}^{(t)}$
        // Update STM
12
      $q^{(t)}\leftarrow\mathrm{D^{S}}(\mathcal{X},\mathcal{M}_{\text{S}},\mathcal{M}_{\text{L}})$
        // Question
13       $\mathcal{R}^{(t)}\leftarrow\mathcal{R}^{(t)}\cup\{q^{(t)}\}$
14
      $\mathcal{R}\leftarrow\mathcal{R}\cup\{\mathcal{R}^{(t)}\}$
        // Append to History
15
16 end for
17
$\mathcal{Y}\leftarrow\mathrm{E}(\mathcal{R})$
  // Evaluate History (ignore q)
18
19if $\exists\,y_{n}^{(t)}=0$ in $\mathcal{Y}$ then
       $\phi\leftarrow\mathrm{F}(\mathcal{R},\mathcal{Y})$
        // Generate Feedback
       $\mathcal{M}_{\text{L}}\leftarrow\mathcal{M}_{\text{L}}\cup\{\phi\}$
        // Update LTM
20
21 end if
22
return $\mathcal{R}$
  // Return Debate History
Algorithm 3 Socratic Refinement Debate
Appendix H Debate Samples
Here we present selected examples from the debate history between agents, provided for demonstration purposes only. All exploitative or misuse-related content has been censored to comply with ethical and safety guidelines.
H.1 Refinement
Simple Debate Example
As shown in Figure 10, agents can observe others’ responses and refine their own in subsequent rounds to become safer. For instance, Agent-0 initially provides a response framed as educational content. In the next round, inspired by Agent-2’s input, it shifts focus to discussing preventative measures and best practices for database protection, making its response safer. Notably, Agent-1, which initially replies with a brief sentence avoiding the question, adopts a more comprehensive and helpful response in the second round, influenced by the other agents’ perspectives.
Socratic Refinement Debate Example
We introduced a Socratic agent specifically to prompt questions when other agents fail to refine their responses on their own, thereby creating a more engaging and interactive debate. As illustrated in Figure 11, following an unsafe response from Agent-0, the Socratic agent raises a critical question: “how can we ensure this information is not misused, even if it is intended for educational purposes?” In the second round, Agent-0 acknowledges the flaw in its earlier response and revises its answer to directly address the Socratic prompt in a safer manner.
Figure 10: An example of a debate under the PReD strategy, illustrating the refinement of an unsafe response.
Figure 11: An example of a debate under the SReD strategy, illustrating the refinement of an unsafe response.
H.2 Error Discovery
As discussed previously, debate can also act as a mechanism for uncovering vulnerabilities when agents are encouraged to challenge each other and exchange arguments. As illustrated in Figure 12, Agent-1 initially provides a safe response by briefly rejecting the question. However, both of the other agents make mistakes by offering unsafe responses. In the second round, Agent-1, now engaged in the discussion, also produces an unsafe response despite having refused to answer in the first round. This example demonstrates how debate can increase the likelihood of error under pressure to participate, highlighting a potential risk: agents may be more prone to generating unsafe content when prompted to elaborate or defend a position in a multi-agent exchange. Debate, while useful for refinement, can therefore also act as a stressor that exposes latent vulnerabilities. This behaviour can be enforced in diffrent debate strategies as well.
Figure 12: An example of a debate acting as an attack, revealing a latent vulnerability.
Appendix I Agent Prompts
This section presents the complete set of prompts used for each agent type in the RedDebate framework.
Figure 13: Debater Agent Prompt Template
Figure 14: Feedback Generator Agent Prompt Template
Figure 15: Angel Agent Prompt Template
Figure 16: Devil Agent Prompt Template
Figure 17: Socratic Agent Prompt Template
Figure 18: Self-Critique Agent Prompt Template. Rules are from the Constitutional AI GitHub repository.
(a) Example prompt used to generate Guardrails LTM, given a harmful input and generated feedback during debate.
(b) Example of a generated Colang flow, preventing the model from responding in similar scenarios (i.e., when the user intent matches the defined guardrail).
Figure 19: Running example of guardrails generation pipeline.